# Layers at Similar Depths Generate Similar Activations Across LLM Architectures

### v1

@@ Abstract
How do the latent spaces used by independently-trained LLMs relate to one another? We study the nearest neighbor relationships induced by activations at different layers of 24 open-weight LLMs, and find that they 1) tend to vary from layer to layer within a model, and 2) are approximately shared between corresponding layers of different models. Claim 2 shows that these nearest neighbor relationships are not arbitrary, as they are shared across models, but Claim 1 shows that they are not “obvious” either, as there is no single set of nearest neighbor relationships that is universally shared. Together, these suggest that LLMs generate a progression of activation geometries from layer to layer, but that this entire progression is largely shared between models, stretched and squeezed to fit into different architectures.
@@ 1 Introduction
At first glance, these nearest neighbors look arbitrary, and it is not clear what relationship they have with the original input or with one another. That makes it all the more surprising that layer 20 of Gemma-2-9B gives exactly the same 3 nearest neighbors (in line with Claim 2). Moreover, layer 30 of Llama-3.1-8B gives another completely different set of nearest neighbors (in line with Claim 1), but many of them are the same as the nearest neighbors given by layer 40 of Gemma-2-9B (Fig. 1).
Llama-3.1-8B and Gemma-2-9B are architecturally different models independently trained by different companies, so it is surprising that we not only find similarities in their latent representations, but that these similarities only appear at corresponding depths. (See Section 2 for a more investigation of this specific example.)
@@ 1.1 Related work
There is an extensive literature on representational similarity in neural networks (Kornblith et al., 2019; Klabunde et al., 2024), and with rising interest in universality in LLMs (Olah et al., 2020; Rai et al., 2025) more recent work has begun to apply representational similarity methods to large transformer models (Huh et al., 2024; Klabunde et al., 2023). However, none have systematically studied affinity matrices comparing layers of architecturally different LLMs or the resulting diagonal structure. Most either 1) compared layers within a single model (Vulić et al., 2020), 2) compared layers between two copies of a model trained with different random seeds (Nguyen et al., 2022; Gurnee et al., 2024), 3) focused on CNNs or very small language models (Raghu et al., 2017), 4) used a representational similarity measure that did not reveal diagonal structure (Lan et al., 2024; Wu et al., 2020; Baek et al., 2024; Brown et al., 2023), or some combination thereof (Kornblith et al., 2019; Nguyen et al., 2021). Of particular note is Wang et al. (2025), who used sparse autoencoders to study representational similarity of Pythia-160M and Mamba-130M and noted that layers at similar depths tended to be more similar to one another.
@@ 2 Motivating experiment
We then repeat this process for layer 20 of Gemma-2-9B, and make a Venn diagram showing which elements of $\mathcal{D}$ were considered to be nearest neighbors to $t$ according to the two layers (Fig. 2).
@@
We now compute the nearest neighbors to $t$ according to layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B (Fig. 3).
@@
That the nearest neighbors according to layer 10 of Llama-3.1-8B and layer 20 of Gemma-2-9B are so similar (as well as layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B) is in line with Claim 2 (that activations from corresponding layers of different models induce similar nearest neighbor relationships). In contrast, that the nearest neighbors according to layer 10 and layer 30 of Llama-3.1-8B are so different is in line with Claim 1 (Fig. 1).
For layer 10 of Llama-3.1-8B and layer 20 of Gemma-2-9B, we identified a superficial property of the inputs that could be dominating the geometry of the latent spaces (i.e., that they mention Twitter). We found no such superficial feature relating nearest neighbors according to layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B, but it is possible that there is one. However, neither of these findings would explain away the main observation: Why should independently trained models identify the same (perhaps superficial) properties to focus on, especially if different layers in the same models focus on different properties?
@@ 3.1 Nearest neighbor relationships
Intuitively, $\mathcal{U}^{(\mathcal{D})}_{k}(x,f)$ gives the $k$ elements of a dataset $\mathcal{D}$ that are closest to $t$, after they have been mapped to vectors by $f$.
@@
The mutual $k$-nearest neighbors representational similarity measure $M^{(\mathcal{D})}_{k}$ is defined by
$M^{(\mathcal{D})}_{k}(f,g)\triangleq\frac{1}{k\|\mathcal{D}\|}\sum_{x\in\mathcal{D}}\left\|\mathcal{U}^{(\mathcal{D})}_{k}(x,f)\cap\mathcal{U}^{(\mathcal{D})}_{k}(x,g)\right\|$
@@
Mutual $k$-NN measures the similarity of the nearest neighbor relationships induced by $f$ to those induced by $g$. For a set of inputs, it gives the expected fraction of the $k$-nearest neighbors that both embedding functions agree on. (Note that $M^{(\mathcal{D})}_{k}$ takes two embedding functions and returns a scalar, and therefore satisfies Definition 2.)
@@ 3.3 Affinity matrices
given some representational similarity measure $s$. Unless otherwise noted, we use $M^{(\mathcal{D})}_{10}$ for $s$, which measures similarity with mutual $k$-NN with $k=10$. The dataset $\mathcal{D}$ must also be specified.
@@
It is worth noting that mutual $k$-NN is always measured with respect to an dataset of inputs $\mathcal{D}$. This extends to affinity matrices generated with mutual $k$-NN as well.
@@ 4 Affinity matrices reveal diagonal structure
Figure 4: Layers at corresponding depths tend to generate similar activations on internet text, as represented by diagonal structure in affinity matrices. 4(a) The affinity matrix of Gemma-2-9B and Llama-3.1-8B on a sample fo texts from OpenWebText. Each cell corresponds to a pair of layers, with one from Gemma-2-9B and the other from Llama-3.1-8B. The value at that cell corresponds to the similarity of nearest neighbor relations induced by the activations at those layers. 4(b) The same as 4(a), but for Llama-3.1-70B and Mistral-7B. 4(c) An array of affinity matrices comparing activations on web text across a wide range of models, all of which show diagonal structure. (See Fig. 12 for an extended version and Fig. 20 for a version that resizes all affinity matrices to squares for comparability.)
@@
We consider 24 open-weight models developed by Google (Gemma Team et al., 2024), Meta (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Microsoft (Abdin et al., 2024), and TII (Almazrouei et al., 2023). (See Appendix C for the full list of models.) For every pair of models, we generate an affinity matrix (Definition 5) which shows how similar the nearest neighbor relationships are between every pair of layers in those models. We measure similarity using mutual $k$-NN (Definition 4) with $k=10$, and using a dataset of 2048 texts randomly sampled from OpenWebText (Gokaslan et al., 2019).
We visualize a selection of the resulting affinity matrices in Fig. 4. The most striking feature of these affinity matrices is their diagonal structure, which is present across a wide range of model comparisons (Fig. 4(c)). This diagonal structure implies that layers at proportionally similar depths generalize more similar activations to one another than other pairs of layers, just as seen in Section 2. This indicates that models generate distinct geometries of activations at different depths, but that activations at a given depth are similar across diverse models. Put another way, language models generate a progression of distinct activation geometries (Claim 1), and this entire progression is largely shared between models (Claim 2).
@@
The most similar layers between models cannot align one-to-one because the models being compared have different numbers of layers. In the case of Llama-3.1-8B and Gemma-2-9B (Fig. 4(a)), the most similar layers are not at the same absolute depths (e.g., layer 10 of each model), but instead at similar depths relative to the total number of layers in the model (e.g., 30% of the way through). This pattern persists across a range of model pairs, and the correspondence between maximally similar layers does not deviate far from proportionality in any of the model pairs we consider (Figs. 5(a) and 5(c)).
@@
When comparing models of vastly different sizes, there is yet more flexibility in the correspondence between layers. For example, consider the affinity matrix comparing Llama-3.1-70B and Mistral-7B, which have 80 and 32 layers respectively (Fig. 4(b)). The early layers appear to correspond almost one-to-one, but the middle layers of Llama-3.1-70B appear highly compressed in Mistral-7B. The later layers are all relatively similar to one another, and so show less diagonal structure. This general pattern is typical, though often stretched and squeezed to fit into models with different numbers of layers (Figs. 4(c) and 12).
Existing work has argued that layers at different depths in transformer language models perform different functions, for example, with early layers dedicated to parsing the input and resolving basic grammar, and later layers dedicated to finding the next token of the response etc. (Tenney et al., 2019; Durrani et al., 2020; Kaplan et al., 2025). It is possible that these functions are reflected in the correspondence between layers of large and small models. Language models need to be able to parse their input, regardless of size, and so the first 15 to 20 layers generally correspond one-to-one with little compression. Middle layers, however, could be responsible for more abstract and optional work that can be highly compressed in smaller models, as is reflected in the affinity matrices.
@@
Intuitively, one might expect there to be higher similarity between models when comparing the earliest and latest layers, as these are the most closely tied to model behavior and the training process. However, we find that the absolute similarity between layers does not vary much with depth, with middle layers being just as similar to one another as early and late layers (Fig. 5(b)). Even in architecturally identical models trained with identical data (but with different random initialization), there is no intrinsic guarantee that that their latent activations be similar, but here we find that similar latent activations are generated at every depth.
@@
The experiments above all look at activations generated on inputs sampled from OpenWebText. In Appendix E, we generate affinity matrices for a range of input datasets, including IMDB movie reviews (Maas et al., 2011), text from books in English and German (Tiedemann, 2012), and various LLM evaluations (Zhou et al., 2023; Hendrycks et al., 2021). The diagonal structure observed with OpenWebText is generally present for other distributions of web text (e.g., IMDB movie reviews). Behavior on other input distributions is investigated in Appendix A.
We also perform experiments with other values of $k$. We find that the choice of $k$ shifts the measured mutual $k$-NN similarity, but does not affect diagonal structure, even in extreme cases like $k=1$ (Appendix D).
Figure 6: 6(a) An affinity matrix (from Fig. 4(a)) overlaid with the generalized diagonal region (red) and the path of maximum similarity (blue). 6(b) Layers pairs that lie on the generalized diagonal have higher mean similarity than those that do not. Each point represents a pair of models, with its position determined by the mean similarity of its layer pairs that lie on and off the generalized diagonal.
@@ 4.1 Statistical test of diagonal structure
For a square matrix $A$, the main diagonal contains those elements $A_{i,j}$ where $i=j$. However, many of the affinity matrices we generate are not square. For a rectangular matrix $A$ with dimensions $n\times m$, we define the generalized diagonal to be those elements $A_{i,j}$ satisfying $j\leq i\leq j+n-m$ (where $n\geq m$ without loss of generality). In the case where $A$ is square (i.e., $n=m$), the generalized diagonal is exactly the usual diagonal. An example of the generalized diagonal is included in Fig. 6(a).
For each affinity matrix, we compute the mean similarity of those pairs of layers that are included in the generalized diagonal, and compare them to the mean similarity of all other pairs (Fig. 6(b)). For every pair of models, we find that the on-diagonal mean similarity is higher than the off-diagonal mean similarity.
For each pair of models, we also perform a one-sided $t$-test and find that for every model pair the on-diagonal mean similarity is greater than the off-diagonal mean similarity with a maximum Bonferroni-adjusted $p$-value of $p=0.0004$ (though most of the $p$-values are much smaller).

### v3

@@ Abstract
How do the latent spaces used by independently-trained LLMs relate to one another? We study the nearest neighbor relationships induced by activations at different layers of 24 open-weight LLMs, and find that they 1) tend to vary from layer to layer within a model, and 2) are approximately shared between corresponding layers of different models. Claim 2 shows that these nearest neighbor relationships are not arbitrary, as they are shared across models, but Claim 1 shows that they are not “obvious” either, as there is no single set of nearest neighbor relationships that is universally shared. Together, these suggest that LLMs generate a progression of activation geometries from layer to layer, but that this entire progression is largely shared between models, stretched and squeezed to fit into different architectures. [^1] Code available at https://github.com/chriswolfram/unireps
@@ 1 Introduction
At first glance, these nearest neighbors look arbitrary, and it is not clear what relationship they have with the original input or with one another. That makes it all the more surprising that layer 20 of Gemma-2-9B gives exactly the same 3 nearest neighbors (in line with Claim 2). Moreover, layer 30 of Llama-3.1-8B gives another completely different set of nearest neighbors (in line with Claim 1), but many of them are the same as the nearest neighbors given by layer 40 of Gemma-2-9B (Figure 1).
Llama-3.1-8B and Gemma-2-9B are architecturally different models independently trained by different companies, so it is surprising that we not only find similarities in their latent representations, but that these similarities only appear at corresponding depths. (See Section 2 for more investigation of this specific example.)
@@ 1.1 Related work
There is an extensive literature on representational similarity in neural networks (Kornblith et al., 2019; Klabunde et al., 2024), and with rising interest in universality in LLMs (Olah et al., 2020; Rai et al., 2025) more recent work has begun to apply representational similarity methods to large transformer models (Huh et al., 2024; Klabunde et al., 2023). However, none have systematically studied affinity matrices comparing layers of architecturally different LLMs or the resulting diagonal structure. Most either 1) compared layers within a single model (Vulić et al., 2020), 2) compared layers between two copies of a model trained with different random seeds (Nguyen et al., 2022; Gurnee et al., 2024), 3) focused on CNNs or very small language models (Raghu et al., 2017), 4) used a representational similarity measure that did not reveal diagonal structure (Lan et al., 2024; Baek et al., 2024; Brown et al., 2023), or some combination thereof (Kornblith et al., 2019; Nguyen et al., 2021). Of particular note is the work of Wang et al. (2025), who used sparse autoencoders to study representational similarity of Pythia-160M and Mamba-130M and noted that layers at similar depths tended to be more similar to one another. Also notable is the work of Wu et al. (2020), who measured the representational similarity of layers from a range of early LLMs. They found that early layers tended to be more similar between models than later layers, and observed diagonal structure in some model comparisons (notably in a comparison of BERT-base and BERT-large).
@@ 2 Motivating experiment
We then repeat this process for layer 20 of Gemma-2-9B, and make a Venn diagram showing which elements of $\mathcal{D}$ were considered to be nearest neighbors to $t$ according to the two layers (Figure 2).
@@
We now compute the nearest neighbors to $t$ according to layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B (Figure 3).
@@
That the nearest neighbors according to layer 10 of Llama-3.1-8B and layer 20 of Gemma-2-9B are so similar (as well as layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B) is in line with Claim 2 (that activations from corresponding layers of different models induce similar nearest neighbor relationships). In contrast, that the nearest neighbors according to layer 10 and layer 30 of Llama-3.1-8B are so different is in line with Claim 1 (Figure 1).
For layer 10 of Llama-3.1-8B and layer 20 of Gemma-2-9B, we identified a superficial property of the inputs that could be dominating the geometry of the latent spaces (i.e., that they mention Twitter). We found no such superficial feature relating nearest neighbors according to layer 30 of Llama-3.1-8B and layer 40 of Gemma-2-9B, but it is possible that there is one. However, neither of these findings would explain away the main observation: Why should independently-trained models identify the same (perhaps superficial) properties to focus on, especially if different layers in the same models focus on different properties?
@@ 3.1 Nearest neighbor relationships
Intuitively, $\mathcal{U}^{{(\mathcal{D})}}_{k}(x,f)$ gives the $k$ elements of a dataset $\mathcal{D}$ that are closest to $x$, after they have been mapped to vectors by $f$.
@@
The mutual $k$-nearest neighbors representational similarity measure $\mathcal{M}^{{(\mathcal{D})}}_{k}$ is defined by
$\mathcal{M}^{{(\mathcal{D})}}_{k}(f,g)\triangleq\frac{1}{k\|\mathcal{D}\|}\sum_{x\in\mathcal{D}}\left\|\mathcal{U}^{{(\mathcal{D})}}_{k}(x,f)\cap\mathcal{U}^{{(\mathcal{D})}}_{k}(x,g)\right\|$
@@
Mutual $k$-NN measures the similarity of the nearest neighbor relationships induced by $f$ to those induced by $g$. For a set of inputs, it gives the expected fraction of the $k$-nearest neighbors that both embedding functions agree on. (Note that $\mathcal{M}^{{(\mathcal{D})}}_{k}$ takes two embedding functions and returns a scalar, and therefore satisfies Definition 2.)
@@ 3.3 Affinity matrices
given some representational similarity measure $s$. Unless otherwise noted, we use $\mathcal{M}^{{(\mathcal{D})}}_{10}$ for $s$, which measures similarity with mutual $k$-NN with $k=10$. The dataset $\mathcal{D}$ must also be specified.
@@
It is worth noting that mutual $k$-NN is always measured with respect to a dataset of inputs $\mathcal{D}$. This extends to affinity matrices generated with mutual $k$-NN as well.
@@ 4 Affinity matrices reveal diagonal structure
Figure 4: Layers at corresponding depths tend to generate similar activations on internet text, as represented by diagonal structure in affinity matrices. 4(a) The affinity matrix of Gemma-2-9B and Llama-3.1-8B on a sample of texts from OpenWebText. Each cell corresponds to a pair of layers, with one from Gemma-2-9B and the other from Llama-3.1-8B. The value at that cell corresponds to the similarity of nearest neighbor relations induced by the activations at those layers. 4(b) The same as 4(a), but for Llama-3.1-70B and Mistral-7B. 4(c) An array of affinity matrices comparing activations on web text across a wide range of models, all of which show diagonal structure. (See Figure 17 for an extended version and Figure 25 for a version that resizes all affinity matrices to squares for comparability.)
@@
We consider 24 open-weight models developed by Google (Gemma Team et al., 2024), Meta (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Microsoft (Abdin et al., 2024), and TII (Almazrouei et al., 2023). (See Appendix F for the full list of models.) For every pair of models, we generate an affinity matrix (Definition 5) which shows how similar the nearest neighbor relationships are between every pair of layers in those models. We measure similarity using mutual $k$-NN (Definition 4) with $k=10$, and using a dataset of 2048 texts randomly sampled from OpenWebText (Gokaslan et al., 2019).
We visualize a selection of the resulting affinity matrices in Figure 4. The most striking feature of these affinity matrices is their diagonal structure, which is present across a wide range of model comparisons (Figure 4(c)). This diagonal structure implies that layers at proportionally similar depths generate more similar activations to one another than other pairs of layers, just as seen in Section 2. This indicates that models generate distinct geometries of activations at different depths, but that activations at a given depth are similar across diverse models. Put another way, language models generate a progression of distinct activation geometries (Claim 1), and this entire progression is largely shared between models (Claim 2).
@@
The most similar layers between models cannot align one-to-one because the models being compared have different numbers of layers. In the case of Llama-3.1-8B and Gemma-2-9B (Figure 4(a)), the most similar layers are not at the same absolute depths (e.g., layer 10 of each model), but instead at similar depths relative to the total number of layers in the model (e.g., 30% of the way through). This pattern persists across a range of model pairs, and the correspondence between maximally similar layers does not deviate far from proportionality in any of the model pairs we consider (Figures 5(a) and 5(c)).
@@
When comparing models of vastly different sizes, there is yet more flexibility in the correspondence between layers. For example, consider the affinity matrix comparing Llama-3.1-70B and Mistral-7B, which have 80 and 32 layers respectively (Figure 4(b)). The early layers appear to correspond almost one-to-one, but the middle layers of Llama-3.1-70B appear highly compressed in Mistral-7B. The later layers are all relatively similar to one another, and so show less diagonal structure. This general pattern is typical, though often stretched and squeezed to fit into models with different numbers of layers (Figures 4(c) and 17).
Existing work has argued that layers at different depths in transformer language models perform different functions, for example, with early layers dedicated to parsing the input and resolving basic grammar, and later layers dedicated to finding the next token of the response etc. (Lad et al., 2024; Tenney et al., 2019; Durrani et al., 2020; Kaplan et al., 2025; Cheng et al., 2025). It is possible that these functions are reflected in the correspondence between layers of large and small models. Language models need to be able to parse their input, regardless of size, and so the first 15 to 20 layers generally correspond one-to-one with little compression. Middle layers, however, could be responsible for more abstract and optional work that can be highly compressed in smaller models, as is reflected in the affinity matrices.
@@
Intuitively, one might expect there to be higher similarity between models when comparing the earliest and latest layers, as these are the most closely tied to model behavior and the training process. Indeed, in their experiments, Wu et al. (2020) found that earlier layers were more similar across models than later layers. However, we find that the absolute similarity between layers does not vary much with depth, with middle layers being just as similar to one another as early and late layers (Figure 5(b)). Even in architecturally identical models trained with identical data (but with different random initialization), there is no intrinsic guarantee that their latent activations be similar, but here we find that similar latent activations are generated at every depth.
@@
The experiments above all look at activations generated on inputs sampled from OpenWebText. In Appendix G, we generate affinity matrices for a range of input datasets, including IMDB movie reviews (Maas et al., 2011), text from books in English and German (Tiedemann, 2012), and various LLM evaluations (Zhou et al., 2023; Hendrycks et al., 2021). The diagonal structure observed with OpenWebText is generally present for other distributions of web text (e.g., IMDB movie reviews). Behavior on other input distributions is investigated in Appendix A.
We also perform experiments with other values of $k$ and find that the choice of $k$ shifts the measured mutual $k$-NN similarity, but does not affect diagonal structure, even in extreme cases like $k=1$ (Appendix D).
Figure 6: 6(a) An affinity matrix (from Figure 4(a)) overlaid with the generalized diagonal region (red). 6(b) Layers pairs that lie on the generalized diagonal have higher mean similarity than those that do not. Each point represents a pair of models, with its position determined by the mean similarity of its layer pairs that lie on and off the generalized diagonal.
@@ 4.1 Statistical test of diagonal structure
Generalized diagonal
For a square matrix $A$, the main diagonal contains those elements $A_{i,j}$ where $i=j$. However, many of the affinity matrices we generate are not square. For a rectangular matrix $A$ with dimensions $n\times m$, we define the generalized diagonal to be those elements $A_{i,j}$ satisfying $j\leq i\leq j+n-m$ (where $n\geq m$ without loss of generality). In the case where $A$ is square (i.e., $n=m$), the generalized diagonal is exactly the usual diagonal. An example of the generalized diagonal is shown in Figure 6(a).
For each affinity matrix (generated with inputs on OpenWebText), we compute the mean similarity of those pairs of layers that are included in the generalized diagonal, and compare them to the mean similarity of all other pairs (Figure 6(b)). For every pair of models, we find that the on-diagonal mean similarity is higher than the off-diagonal mean similarity.
Naive $t$-test
For each model pair $(M_{u},M_{v})$ we then have the null hypothesis $H_{0}^{u,v}:\mu_{u,v}\leq\mu^{\prime}_{u,v}$ where $\mu_{u,v}$ denotes the mean similarity of on-diagonal cells in the affinity matrix comparing $M_{u}$ and $M_{v}$, and $\mu^{\prime}_{u,v}$ denotes the mean similarity of off-diagonal cells. A naive approach to generate a $p$-value for $H_{0}^{u,v}$, which assumes that the affinity matrix’s entries are independent, is to perform a one-sided $t$-test comparing on- and off-diagonal cells in the affinity matrix for $M_{u}$ and $M_{v}$. Performing such a $t$-test for each model pair, we find that the resulting $p$-values are astronomically small, with a maximum of $8\times 10^{-7}$. Moreover, the $p$-value for the union null hypothesis
$H_{0}=\bigcup_{u,v}H_{0}^{u,v}:\exists_{u,v}~\mu_{u,v}\leq\mu^{\prime}_{u,v}$
is given by the maximum of all of the constituent $p$-values (Berger, 1982; Kim & Allison, 2006). Therefore, according to the naive $t$-tests, the $p$-value for $H_{0}$ is $8\times 10^{-7}$, indicating that every pair of models exhibit statistically higher similarity on the diagonal than off of it.
Block bootstrap
The $t$-test approach assumes independence of the cells in each affinity matrix. However, this assumption is clearly violated, as any pair of cells in a row ii—e.g., $A^{u,v}_{i,j}$ and $A^{u,v}_{i,j^{\prime}}$—for the affinity matrix comparing models $M_{u}$ and $M_{v}$ are both functions of the same embedding function $f^{{(i)}}_{M_{u}}$. The same is true for pairs of cells in any column $j$, both of which are functions of $f^{{(j)}}_{M_{v}}$. In addition, the embedding functions at nearby layers within a model—e.g., $f^{{(i)}}_{M_{u}}$ and $f^{{(i+1)}}_{M_{u}}$—may be dependent due to residual connections and other aspects of the transformer architecture.
To account for these local dependencies, we use the moving block bootstrap (Lahiri, 1999; Davison & Hinkley, 1999) with a block size of $5\times 5$ to sample from a null distribution which preserves local dependencies but scrambles diagonal structure (see Appendix B for more details). By comparing samples from this null distribution to the actual affinity matrix, we can estimate the probability that the difference of on- and off-diagonal means would be as great under the null model as those of the observed affinity matrix. Repeating this process for every affinity matrix, we find a maximum $p$-value of $4.84\times 10^{-3}$, again indicating that the on-diagonal mean is greater than the off-diagonal mean for every pair of models.